# Bridging Draft Policy Misalignment: Group Tree Optimization for Speculative Decoding

### v1

@@ Abstract
Speculative decoding accelerates large language model (LLM) inference by letting a lightweight draft model propose multiple tokens that the target model verifies in parallel. Yet existing training objectives optimize only a single greedy draft path, while decoding follows a tree policy that re-ranks and verifies multiple branches. This draft policy misalignment limits achievable speedups. We introduce Group Tree Optimization (GTO), which aligns training with the decoding-time tree policy through two components: (i) Draft Tree Reward, a sampling-free objective equal to the expected acceptance length of the draft tree under the target model, directly measuring decoding performance; (ii) Group-based Draft Policy Training, a stable optimization scheme that contrasts trees from the current and a frozen reference draft model, forming debiased group-standardized advantages and applying a PPO-style surrogate along the longest accepted sequence for robust updates. We further prove that increasing our Draft Tree Reward provably improves acceptance length and speedup. Across dialogue (MT-Bench), code (HumanEval), and math (GSM8K), and multiple LLMs (e.g., LLaMA-3.1-8B, LLaMA-3.3-70B, Vicuna-1.3-13B, DeepSeek-R1-Distill-LLaMA-8B), GTO increases acceptance length by $7.4\%$ and yields an additional $7.7\%$ speedup over prior state-of-the-art EAGLE-3. By bridging draft policy misalignment, GTO offers a practical, general solution for efficient LLM inference.
@@ 4 Experiment
Models & datasets. We test GTO on a representative set of LLMs, including LLaMA-3.1-Instruct-8B (Touvron et al., 2023b), LLaMA-3.3-Instruct-70B (Touvron et al., 2023b), Vicuna-1.3-13B (Fan et al., 2025), and DeepSeek-R1-Distill-LLaMA-8B (Guo et al., 2025). All experiments are conducted on a single NVIDIA A100 80GB GPU, except for LLaMA-3.3-70B, which requires two GPUs. We benchmark performance on three widely used evaluation suites: multi-turn conversation (MT-Bench (Zheng et al., 2023)), code generation (HumanEval (Chen et al., 2021)), and mathematical reasoning (GSM8K (Cobbe et al., 2021)).
@@
By default, GTO initializes its draft model from the one provided by EAGLE-3. To assess compatibility, we also experiment with draft models trained by other approaches (see Table 2). The initialized draft models are then fine-tuned with GTO on the ShareGPT dataset (Chiang et al., 2023), except for the reasoning model DeepSeek-R1-Distill-LLaMA 8B, which is fine-tuned on OpenThoughts-114k-math dataset (Guha et al., 2025). See additional training details for GTO in Appendix B, and details for the baselines in Appendix C.
@@
LLaMA-3.1
Instruct
8B
PLD 1.53 1.61 1.69 1.73 1.79 1.85 1.67 1.73 N/A, since the acceptance conditions are relaxed
@@
DeepSeek-R1
Distill-LLaMA
8B
PLD 1.34 1.42 1.53 1.62 1.48 1.54 1.45 1.53 N/A, since the acceptance conditions are relaxed
@@
LLaMA-3.3
Instruct
70B
PLD 1.43 1.51 1.58 1.67 1.52 1.61 1.51 1.60 N/A, since the acceptance conditions are relaxed
@@ 4.1 Main Results
LLaMA-3
Instruct
8B
GRIFFIN 3.09 4.85 3.65 5.97 3.30 5.31 3.35 5.38 2.62 4.35 3.31 5.62 3.07 5.08 3.00 5.02
@@
LLaMA-2
Chat
7B
GRIFFIN 3.12 5.11 3.61 5.93 3.10 5.27 3.28 5.44 2.81 4.81 3.33 5.63 3.06 5.26 3.07 5.23
@@ Appendix D Training Overhead of GTO
All results were obtained on NVIDIA A100 80 GB GPUs under mixed precision with ZeRO-2. The Phase-II GTO fine-tuning requires approximately (i) 200 GPU-hours for 7B models, (ii) 400 GPU-hours for 13B models, and (iii) 900 GPU-hours for 70B models. These compute budgets cover end-to-end GTO training (including grouped tree construction and verification) and exclude any pretraining of the base or drafter models, as we fine-tune on publicly available pretrained drafters.

### v2

# arXiv:2509.22134v2[cs.CL] 28 Feb 2026

@@ Abstract
Speculative decoding accelerates large language model (LLM) inference by letting a lightweight draft model propose multiple tokens that the target model verifies in parallel. Yet existing training objectives optimize only a single greedy draft path, while decoding follows a tree policy that re-ranks and verifies multiple branches. This draft policy misalignment limits achievable speedups. We introduce Group Tree Optimization (GTO), which aligns training with the decoding-time tree policy through two components: (i) Draft Tree Reward, a sampling-free objective equal to the expected acceptance length of the draft tree under the target model, directly measuring decoding performance; (ii) Group-based Draft Policy Training, a stable optimization scheme that contrasts trees from the current and a frozen reference draft model, forming debiased group-standardized advantages and applying a PPO-style surrogate along the longest accepted sequence for robust updates. We further prove that increasing our Draft Tree Reward provably improves acceptance length and speedup. Across dialogue (MT-Bench), code (HumanEval), and math (GSM8K), and multiple LLMs (e.g., LLaMA-3.1-8B, LLaMA-3.3-70B, Vicuna-1.3-13B, DeepSeek-R1-Distill-LLaMA-8B, Qwen3-8B), GTO increases acceptance length by $7.4\%$ and yields an additional $7.7\%$ speedup over prior state-of-the-art EAGLE-3. By bridging draft policy misalignment, GTO offers a practical, general solution for efficient LLM inference. Code and draft models are available at https://github.com/hsj576/GTO.
@@ 4 Experiment
Models & datasets. We test GTO on a representative set of LLMs, including LLaMA-3.1-Instruct-8B (Touvron et al., 2023b), LLaMA-3.3-Instruct-70B (Touvron et al., 2023b), Vicuna-1.3-13B (Fan et al., 2025), DeepSeek-R1-Distill-LLaMA-8B (Guo et al., 2025) and Qwen3-8B (Yang et al., 2025). All experiments are conducted on a single NVIDIA A100 80GB GPU, except for LLaMA-3.3-70B, which requires two GPUs. We benchmark performance on three widely used evaluation suites: multi-turn conversation (MT-Bench (Zheng et al., 2023)), code generation (HumanEval (Chen et al., 2021)), and mathematical reasoning (GSM8K (Cobbe et al., 2021)).
@@
By default, GTO initializes its draft model from the one provided by EAGLE-3. To assess compatibility, we also experiment with draft models trained by other approaches (see Table 2). The initialized draft models are then fine-tuned with GTO on the ShareGPT dataset (Chiang et al., 2023), except for the reasoning model DeepSeek-R1-Distill-LLaMA 8B, which is also fine-tuned on OpenThoughts-114k-math dataset (Guha et al., 2025). See additional training details for GTO in Appendix B, and details for the baselines in Appendix C.
@@
LLaMA-3.1 Instruct 8B PLD 1.53 1.61 1.69 1.73 1.79 1.85 1.67 1.73 N/A, since the acceptance conditions are relaxed
@@
DeepSeek-R1 Distill-LLaMA 8B PLD 1.34 1.42 1.53 1.62 1.48 1.54 1.45 1.53 N/A, since the acceptance conditions are relaxed
@@
LLaMA-3.3 Instruct 70B PLD 1.43 1.51 1.58 1.67 1.52 1.61 1.51 1.60 N/A, since the acceptance conditions are relaxed
@@ 4.1 Main Results
LLaMA-3 Instruct 8B GRIFFIN 3.09 4.85 3.65 5.97 3.30 5.31 3.35 5.38 2.62 4.35 3.31 5.62 3.07 5.08 3.00 5.02
@@
LLaMA-2 Chat 7B GRIFFIN 3.12 5.11 3.61 5.93 3.10 5.27 3.28 5.44 2.81 4.81 3.33 5.63 3.06 5.26 3.07 5.23
@@ 4.2 Ablation Study
These observations align with the theoretical insights from the GRPO (Shao et al., 2024), which shows that very small groups suffer from high-variance and unstable updates, while very large groups suffer from signal attenuation and slower convergence due to excessive averaging. In our GTO experiments, we also observe this qualitative pattern. Group sizes in the range of 4–8 consistently provide the best balance: they significantly reduce variance relative to size 1, yet still preserve enough reward contrast to produce strong learning signals. Larger groups (e.g., size 32) remain stable but deliver noticeably weaker improvements due to the normalization-induced compression described above.
@@
Table 6: Ablation of continual training on draft model.
Temperature = 0 Temperature = 1
continual training 3.37 5.38 3.61 5.94 3.81 6.24 3.60 5.85 2.70 4.53 3.16 5.15 3.54 5.93 3.13 5.20
Continual training.
To ensure that the performance gains of GTO stem from our proposed algorithmic improvements rather than the additional computational budget, we introduce a continual training baseline. Specifically, we further fine-tune the vanilla EAGLE-3 draft model for an additional 200 A100-80G GPU hours. This baseline employs the exact same training data and strictly follows the original EAGLE-3 training recipe for the DeepSeek-R1-Distill-LLaMA-8B target model.
As reported in Table 6, GTO achieves a consistently higher speedup (approximately 4%–7% on average) across various benchmarks and temperatures compared to this continued-training baseline. Furthermore, we observe that the continual training baseline performs almost identically to the vanilla EAGLE-3. This behavior is expected: because the training dataset is already encompassed within EAGLE-3’s original training mixture, the draft model is already near convergence on this distribution. Without the introduction of novel data, allocating additional compute to the standard training objective merely reinforces an already-optimized model, yielding negligible improvements.
These findings confirm that the efficacy of GTO does not derive from simply extending training time or data exposure. Instead, the improvements are fundamentally algorithmic, rooted in GTO’s ability to explicitly align the draft tree policy with the target acceptance dynamics, thereby directly resolving the policy misalignment that typically bottlenecks speculative decoding.
@@ 5 Conclusion
Acknowledgement
This work was supported by the Yangtze River Delta Science and Technology Innovation Community Joint Research Project (YDZX20233100004031) and the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (Proposal ID: 25-SIS-SMU-003). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
Ethics Statement
GTO improves efficiency of large language model decoding. Nevertheless, faster generation could increase the throughput of undesirable content if deployed without safeguards. We recommend deploying GTO only with established safety measures (content filters, rate limiting, audit logging, and red-teaming) and within the original safety and usage policies of the underlying models.
@@ B.3 Training Configuration
Additional hyperparameters and scripts are available at https://github.com/hsj576/GTO.
@@ Appendix D Training Overhead of GTO
All results were obtained on NVIDIA A100 80 GB GPUs under mixed precision with ZeRO-2. The Phase-II GTO fine-tuning requires approximately (i) 200 GPU-hours for 8B models, (ii) 400 GPU-hours for 13B models, and (iii) 900 GPU-hours for 70B models. These compute budgets cover end-to-end GTO training (including grouped tree construction and verification) and exclude any pretraining of the base or drafter models, as we fine-tune on publicly available pretrained drafters.
@@
Appendix E Ablation on Tree Configuration
In our primary experiments, we default to the standard EAGLE tree configuration (e.g., depth 7, 60 total tokens). This tree structure has emerged as the de facto standard in modern speculative decoding research and is widely adopted by recent state-of-the-art frameworks (e.g., EAGLE-3, HASS, GRIFFIN). By deliberately adopting this configuration, we ensure that our results are strictly comparable to established baselines, accurately reflect realistic deployment settings, and maximize practical relevance for the community.
Tree-Agnostic Formulation.
Importantly, GTO is not inherently constrained to any specific tree layout. Methodologically, GTO optimizes the drafter to maximize the expected acceptance length under speculative decoding. Because the training procedure learns underlying acceptance-related token distribution patterns rather than memorizing a fixed tree shape, GTO is conceptually compatible with any tree configuration that satisfies two basic conditions: (i) the drafter can successfully construct the tree, and (ii) a tree-level reward can be computed via target model verification. This design ensures that GTO’s optimization principles generalize well beyond the default EAGLE layout.
Interpolation and Extrapolation Performance.
To empirically validate GTO’s robustness to varying tree structures, we conduct an ablation study exploring the extrapolation and interpolation of the tree configuration at inference time. Specifically, we take a drafter trained exclusively on the standard EAGLE tree (depth 7, 60 total tokens) and evaluate it using modified configurations during inference. We vary the tree depth $\in\{6,7,8\}$ and the total drafted tokens per step $\in\{50,60,70\}$. The comprehensive results across different benchmarks and temperatures are presented in Table 7. We observe four key findings:
  • Consistent Performance: GTO maintains strong speedup ($SR$) and acceptance rate ($\tau$) improvements across all 9 tested configurations, demonstrating high robustness to both depth and width variations.
  • Graceful Interpolation: Configurations closely interpolating the training setting exhibit minimal performance degradation, with speedup varying by only $1\%$$3\%$ from the optimal setting.
  • Effective Extrapolation: When extrapolating to depth 8 (beyond the training depth of 7), GTO still achieves competitive or even superior performance. This indicates that the token-level alignment patterns learned by GTO transfer effectively to deeper and larger trees.
  • Stability Across Conditions: The robustness holds consistently across MT-bench, HumanEval, and GSM8K, as well as under both greedy ($T=0$) and sampling ($T>0$) decoding settings. This confirms that the generalization is neither task-specific nor decoding-strategy-specific.
In conclusion, these results demonstrate that GTO does not overfit to a specific tree configuration during the training phase. The learned alignment strategies remain highly effective when the tree structure is modified at inference time. Consequently, GTO fully supports dynamic tree configuration updates during inference while maintaining, and in some cases even improving, overall speculative decoding performance.
Table 7: Ablation of Tree Configuration $m$ on LLaMA-3.1 8B.
Temperature = 0 Temperature = 1
Depth Total Tokens MT-bench HumanEval GSM8K Average MT-bench HumanEval GSM8K Average
6 50 3.29 5.81 3.71 6.45 3.46 6.12 3.49 6.13 2.39 4.54 3.10 5.75 2.90 5.40 2.80 5.23
6 60 3.32 5.86 3.73 6.50 3.49 6.21 3.51 6.19 2.41 4.56 3.12 5.79 2.93 5.51 2.82 5.29
6 70 3.34 5.93 3.73 6.52 3.51 6.27 3.53 6.24 2.41 4.56 3.12 5.80 2.93 5.52 2.82 5.29
7 50 3.39 6.05 3.95 6.87 3.56 6.39 3.63 6.44 2.44 4.60 3.15 5.86 2.83 5.30 2.80 5.25
7 60 3.44 6.15 4.17 6.95 3.59 6.47 3.73 6.52 2.49 4.70 3.17 5.92 3.02 5.75 2.89 5.46
7 70 3.46 6.23 4.17 7.00 3.63 6.57 3.76 6.60 2.51 4.74 3.24 6.10 3.10 5.86 2.95 5.57
8 50 3.56 6.22 4.22 7.24 3.61 6.59 3.80 6.68 2.49 4.72 3.29 6.34 2.80 5.24 2.86 5.43
8 60 3.54 6.34 4.24 7.33 3.66 6.69 3.81 6.79 2.46 4.73 3.20 6.12 2.95 5.56 2.87 5.47
8 70 3.49 6.43 4.20 7.39 3.71 6.83 3.80 6.88 2.46 4.74 3.22 6.28 2.78 5.26 2.82 5.43
Appendix F LLM Usage Statement
Large language models were used minimally for proofreading and grammar checking. The research ideas, methodology, experiments, and analysis were entirely conceived and conducted by the authors.